# Faster Diffusion Models via Higher-Order Approximation

### v1

@@ Abstract
score function evaluations (up to log factor) in the presence of accurate scores, where $K$ is an arbitrarily large fixed integer. This result applies to a broad class of target data distributions, without the need for assumptions such as smoothness or log-concavity. Our theory is robust vis-a-vis inexact score estimation, degrading gracefully as the score estimation error increases — without demanding higher-order smoothness on the score estimates as assumed in previous work. The proposed algorithm draws insight from high-order ODE solvers, leveraging high-order Lagrange interpolation and successive refinement to approximate the integral derived from the probability flow ODE.
@@ 1.1 Diffusion models and score function evaluations
(forward $(\text{forward process})\qquad X_{0}$ $\sim p_{\mathsf{data}},\quad X_{t}=\sqrt{1-\beta_{t}}\,X_{t-1}+\sqrt{\beta_{t}}\,W_{t},\quad t=1,\cdots,T,$ (1a)
@@
(backward $(\text{backward process})\qquad Y_{T}$ $\sim\mathcal{N}(0,I_{d}),\qquad Y_{T}\rightarrow Y_{T-1}\rightarrow\dots\rightarrow Y_{1}\rightarrow Y_{0}$
@@ 1.2 Training-free acceleration
In this work, we explore principled, training-free acceleration of diffusion models (Lu et al.,, 2022; Zheng et al.,, 2023; Zhao et al.,, 2024; Zhang and Chen,, 2023; Jolicoeur-Martineau et al.,, 2021; Xue et al.,, 2024). In a nutshell, the training-free approach wraps around and makes more efficient use of the pretrained score functions to expedite data generation, which stands in sharp contrast to the training-based approach — such as distillation or consistency models — that performs another round of resource-intensive training (Luhman and Luhman,, 2021; Salimans and Ho,, 2022; Meng et al.,, 2023; Song et al.,, 2023; Li et al., 2024b, ; Nichol and Dhariwal,, 2021). Remarkably, training-free acceleration has already seen widespread adoption in practice, with a prominent example being the DPM-solver (Luhman and Luhman,, 2021) that has achieved dramatic empirical speedup compared to the original probability flow ODE algorithm.
Nevertheless, the theoretical underpinnings for accelerated diffusion models remain largely elusive. It was not until recently that researchers began developing non-asymptotic analysis frameworks for the training-free approach, although significant technical hurdles remain that limit the theoretical performance. More concretely, consider training-free acceleration methods that come with provable guarantees, and suppose that we would like to generate a sample whose distribution is $\varepsilon$-close in total-variation (TV) distance to the target distribution of interest. Li et al., 2024a proposed a deterministic accelerated approach, akin to the second-order ODE solver, achieving an iteration complexity of $\widetilde{O}(d^{3}/\sqrt{\varepsilon})$. [^1] Throughout this paper, the standard notation $f(d,\varepsilon^{-1})=O(g(d,\varepsilon^{-1}))$ or $f(d,\varepsilon^{-1})\lesssim g(d,\varepsilon^{-1})$ means that there exists a numerical constant $c_{1}>0$ such that $|f(d,\varepsilon^{-1})|\leq c_{1}|g(d,\varepsilon^{-1})|$; $f(d,\varepsilon^{-1})\gtrsim g(d,\varepsilon^{-1})$ means that $g(d,\varepsilon^{-1})\lesssim f(d,\varepsilon^{-1})$; and $f(d,\varepsilon^{-1})\asymp g(d,\varepsilon^{-1})$ means that both $f(d,\varepsilon^{-1})\lesssim g(d,\varepsilon^{-1})$ and $f(d,\varepsilon^{-1})\gtrsim g(d,\varepsilon^{-1})$ hold true. The notation $\widetilde{O}(\cdot)$ is defined analogously to $O(\cdot)$ except that the log dependency is hidden. Huang et al., 2024a analyzed the $p$-th order Runge-Kutta method for solving the probability flow ODE, and established an iteration complexity of $\widetilde{O}\big{(}(LDd)^{1+1/p}/\varepsilon^{1/p}\big{)}$ for any constant order $p$, where $D$ represents the radius of the data support and $L$ bounds certain high-order derivatives of the score estimates. This represents a clear improvement over Li et al., 2024a in terms of the $\varepsilon$ dependency, although the stringent requirements on $L$ and $D$ limit the range of data distributions it can accommodate without compromising the acceleration effect. Meanwhile, stochastic training-free methods have been proposed and analyzed in recent literature as well (Li et al., 2024a, ; Wu et al.,, 2024; Li and Cai,, 2024), with the state-of-the-art results achieving an iteration complexity of $\widetilde{O}(d^{5/4}/\sqrt{\varepsilon})$ (Li and Cai,, 2024). While such stochastic acceleration methods often come with theoretical guarantees under minimal assumptions, their $\varepsilon$ dependency remains considerably worse than that of the deterministic counterpart.
@@ 1.3 Main contributions
in the presence of accurate scores. Our theory does not require any smoothness or log-concavity assumptions on the target distribution, making it broadly applicable. When only inexact score estimates are available, we demonstrate that the sampling quality degrades gracefully as the score estimation error and the associated Jacobian errors increase. Our results outperform Li et al., 2024a ; Wu et al., (2024); Li and Cai, (2024) by offering better scaling in both $\varepsilon$ and $d$, and improve upon Huang et al., 2024a by accommodating a broader set of nonsmooth distributions without assuming higher-order smoothness.
@@ 1.4 Other related work
The development of convergence theory for diffusion models — particularly DDPM and DDIM — has received much attention during the past few years. Partial examples include Chen et al., (2022); Liu et al., (2022); Lee et al., (2023); Chen et al., 2023a ; Chen et al., 2023c ; Li et al., (2023); Cheng et al., (2023); Chen et al., 2023b ; Benton et al., (2024); Tang and Zhao, (2024); Liang et al., (2024); Li et al., (2025); Jiao and Li, (2024); Cai and Li, (2025); Li et al., 2024d ; Gao and Zhu, (2024); Li and Jiao, (2024); Li and Yan, 2024b ; see also the references therein. The convergence analysis for the DDPM was carried out in Chen et al., (2022) by means of Girsanov’s theorem, and this analysis framework was subsequently improved to accommodate nonsmooth data distributions Lee et al., (2023); Chen et al., 2023a , unveiling an iteration complexity with nearly linear $d$ dependency Benton et al., (2024); Li and Yan, 2024b . When it comes to the DDIM sampler or the probability flow ODE, the convergence analysis was originally provided in Chen et al., 2023c , with a set of subsequent work devoted to sharpening the iteration complexity (Li et al., 2024c, ; Huang et al., 2024a, ; Li et al., 2024d, ; Liang et al.,, 2025). In addition, a recent line of work Li and Yan, 2024a ; Huang et al., 2024b ; Azangulov et al., (2024); Potaptchik et al., (2024); Liang et al., (2025); Tang and Yan, (2025) explored how DDPM and DDIM adapt to unknown low-dimensional structures underlying the target data distribution, leading to substantially improved theoretical guarantees.
@@ 2 Preliminaries
where $X_{0}$ and $Z$ are independently generated. Clearly, with this construction we have
@@
In the sequel, we will often refer to $\tau$ as the time variable too as long as it is clear from the context, and will develop and describe our algorithm mainly based on this continuous-time forward process (4).
@@
for any $x\in\mathbb{R}^{d}$, which implies that
@@
As has been shown by Song et al., (2021), there exists an ordinary differential equation (ODE), called the probability flow ODE, that is able to reverse the forward process (4) (in the sense of yielding matching marginal distributions). Note that the precise form of the probability flow ODE depends on the parameterization of the forward process (e.g., $(X_{t})$ and $(\overline{X}_{\tau})$ are associated with different, although related, ODEs). In particular, the probability flow ODE associated with process (4) takes the following form:
@@ 3 Algorithm
In this section, we present the main ideas and detailed procedure of the proposed algorithm. At a high level, our algorithm is an iterative procedure that attempts to approximately solve ODE (9) by approximating the true scores in the integrand using high-order polynomials.
@@
Recall that $X_{t}=\overline{X}_{1-\overline{\alpha}_{t}}$ and $X_{t-1}=\overline{X}_{1-\overline{\alpha}_{t-1}},$ which makes apparent the importance of time points $1-\overline{\alpha}_{t}$ and $1-\overline{\alpha}_{t-1}$ when generating $Y_{t-1}$. Armed with the probability flow ODE (9), an ideal strategy for step $t-1$ would be to evaluate the following integral:
@@
However, computing (10) involves evaluation an infinite number of score functions (i.e., those in the integral), making it computationally intensive. As a result, everything comes down to how to approximate the integral using a small number of score function evaluations.
1 input: $\tau_{t,0}=\tau_{t+1,K-1}=1-\overline{\alpha}_{t}$ for all $11\leq t\leq T-$ 1, and $\tau_{T,0}=1-\overline{\alpha}_{T}$. Choose $\tau_{t,i}$ as in (11).
2 initialization: $t=T,\dots,2$, $x_{\tau_{T,0}}\leftarrow Y_{T}\sim\mathcal{N}(0,I_{d})$.
3 for $t>1$ do
@@ 7      $Y_{t-1}\leftarrow x_{\tau_{t,K-1}}^{(N)}$
Higher-order polynomials.
In comparison to first-order solvers like DDIM that construct $Y_{t-1}$ solely based on the score $s_{1-\overline{\alpha}_{t}}(\cdot)$, we propose to take advantage of the score functions at $K$ time points — denoted by $\tau_{t,0},\dots,\tau_{t,K-1}$ — within the interval $[1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$. To be precise, set
@@
to be the two endpoints of $[1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$, and create equi-spaced points within this interval as
@@
$\psi_{i}(\tau)\coloneqq\frac{\prod_{i^{\prime}:i^{\prime}\neq i}(\tau-\tau_{t,i^{\prime}})}{\prod_{i^{\prime}:i^{\prime}\neq i}(\tau_{t,i}-\tau_{t,i^{\prime}})},\qquad 0\leq i\leq K.$ (12)
Given a set of points $(\tau_{t,i},(1-\tau_{t,i})^{-3/2}s_{\tau_{t,i}}(Y^{\mathsf{ode}}_{\tau_{t,i}}))$ for $10\leq i\leq K-$ 1, we would like to approximate $\frac{1}{(1-\tau)^{3/2}}s^{\star}_{\tau}(x_{\tau})$ via the Lagrange interpolating polynomial passing through the above $K$ points, i.e.,
@@
Here, we have also replaced the true score $s_{t}^{\star}$ with the score estimate $s_{t}$ on the right-hand side of (13). Note that the Lagrange interpolating polynomial on right-hand side of (13) forms a unique degree-$(K-1)$ polynomial passing through these $K$ points.
@@
Noteworthily, the polynomial appproximation (13) still cannot be readily used, given that the points $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$ are inaccessible in general. To remedy this issue, we propose a successive refinement scheme that alternates between estimating $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$ and computing the the associated scores. More precisely, let us perform an iterative estimation procedure containing $N$ rounds, where in each round $n$, we produce a sequence $\{x_{\tau_{t},i}^{(n)}\}_{0\leq i\leq K-1}$ as an estimate of $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$. When $\{x_{\tau_{t},i}^{(n)}\}_{0\leq i\leq K-1}$ is available, we compute, for each $\tau^{\prime}\in\{\tau_{t,0},\dots,\tau_{t,K-1}\}$,
@@
In a nutshell, we employ $\{x_{\tau_{t},i}^{(n)}\}_{0\leq i\leq K-1}$ to update the scores used in the polynomial approximation (13), which in turn helps us generate $\{x_{\tau_{t},i}^{(n+1)}\}_{0\leq i\leq K-1}$ as improved estimates of $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$.
@@
We are now ready to present our algorithm.
@@
  • Iterative update rule. Working backward from $t=T$ to $t=1$, we generate the point $Y_{t-1}$ as follows. Let us look at the time points $\{\tau_{t,i}\}_{0\leq i<K}$, and initialize both the sequence $\{x_{\tau_{t,i}}^{(0)}\}_{0\leq i\leq K-1}$ and $x_{\tau_{t,0}}$ using the value $x_{\tau_{t+1,K-1}}^{(N)}$ obtained in the previous iteration. For each round $n=0,\dots,N-1$, we generate a sequence $\{x_{\tau_{t,i}}^{(n+1)}\}_{0\leq i\leq K-1}$ according to Eq. (14). After $N$ rounds of successive refinement, we output $Y_{t-1}=x_{\tau_{t,K-1}}^{(N)}$.
@@
Several remarks are in order. To begin with, once the value of $Y_{T}$ is given, the sequence $\{x_{\tau_{t,i}}^{(n)}\}$ is purely deterministic. Thus, when accurate score functions are available at the requested time points, the discrepancy between the distributions of $Y_{t}$ and $X_{t}$ mainly stems from the approximation error when interpolating a complicated function using polynomials of degree $K-1$. In addition, our algorithm requires computing the score functions at the sequence $\{x_{\tau_{t,i}}^{(n)}\}_{0\leq i\leq K-1,0\leq n\leq N,1\leq t\leq T}$, and hence the total number of score function evaluations is $TKN$. As we shall see momentarily in our theoretical development, $K$ is taken to be a constant, whereas $N$ is chosen to be logarithmically large. Consequently, the total number of score function evaluations is on the order of $\widetilde{O}(T)$.

### v2

@@ Abstract
score function evaluations (up to log factor) in the presence of accurate scores, where $K>0$ is an arbitrary fixed integer. This result applies to a broad class of target data distributions, without the need for assumptions such as smoothness or log-concavity. Our theory is robust vis-a-vis inexact score estimation, degrading gracefully as the score estimation error increases — without demanding higher-order smoothness on the score estimates as assumed in previous work. The proposed algorithm draws insight from high-order ODE solvers, leveraging high-order Lagrange interpolation and successive refinement to approximate the integral derived from the probability flow ODE. More broadly, our work develops a theoretical framework towards understanding the efficacy of high-order methods for accelerated sampling.
Keywords: diffusion models, probability flow ODE, high-order ODE, sampling
@@ 1.1 Diffusion models and score function evaluations
$(\text{forward process})\qquad X_{0}$ $\sim p_{\mathsf{data}},\quad X_{t}=\sqrt{1-\beta_{t}}\,X_{t-1}+\sqrt{\beta_{t}}\,W_{t},\quad t=1,\cdots,T,$ (1a)
@@
$(\text{backward process})\qquad Y_{T}$ $\sim\mathcal{N}(0,I_{d}),\qquad Y_{T}\rightarrow Y_{T-1}\rightarrow\dots\rightarrow Y_{1}\rightarrow Y_{0}$
@@ 1.2 Training-free acceleration
In this work, we explore principled, training-free acceleration of diffusion models (Lu et al.,, 2022; Zheng et al.,, 2023; Zhao et al.,, 2024; Zhang and Chen,, 2023; Jolicoeur-Martineau et al.,, 2021; Xue et al.,, 2024). In a nutshell, the training-free approach wraps around and makes more efficient use of the pretrained score functions to expedite data generation, which stands in sharp contrast to the training-based approach — such as distillation or consistency models — that performs another round of resource-intensive training (Luhman and Luhman,, 2021; Salimans and Ho,, 2022; Meng et al.,, 2023; Song et al.,, 2023; Li et al., 2024b, ; Nichol and Dhariwal,, 2021). Remarkably, training-free acceleration has already seen widespread adoption in practice, with prominent examples being the DPM-solver (Luhman and Luhman,, 2021), DPM-Solver++ (Lu et al.,, 2025), and UNiPC (Zhao et al.,, 2023) that have achieved dramatic empirical speedups compared to the original probability flow ODE algorithm.
[…]
@@ 1.3 Main contributions
in the presence of accurate scores. Our theory does not require any smoothness or log-concavity assumptions on the target distribution, making it broadly applicable. When only inexact score estimates are available, we demonstrate that the sampling quality degrades gracefully as the score estimation error and the associated Jacobian errors increase. Our results outperform Li et al., 2024a ; Wu et al., (2024); Li and Cai, (2024) by offering better scaling in both $\varepsilon$ and $d$, and improve upon Huang et al., 2024a by accommodating a broader set of nonsmooth distributions without assuming higher-order smoothness. More broadly, our theory develops a theoretical framework that proves effective for analyzing higher-order accelerated samplers.
@@ 1.4 Other related work
The development of convergence theory for diffusion models — particularly DDPM and DDIM — has received much attention during the past few years. Partial examples include Chen et al., (2022); Liu et al., (2022); Lee et al., (2023); Chen et al., 2023a ; Chen et al., 2023c ; Li et al., (2023); Cheng et al., (2023); Chen et al., 2023b ; Benton et al., (2024); Tang and Zhao, (2024); Liang et al., (2024); Li et al., (2025); Jiao and Li, (2024); Cai and Li, (2025); Li et al., 2024d ; Gao and Zhu, (2024); Li and Jiao, (2024); Li and Yan, 2024b ; see also the references therein. The convergence analysis for the DDPM was carried out in Chen et al., (2022) by means of Girsanov’s theorem, and this analysis framework was subsequently improved to accommodate nonsmooth data distributions Lee et al., (2023); Chen et al., 2023a , unveiling an iteration complexity with nearly linear $d$ dependency Benton et al., (2024); Li and Yan, 2024b . When it comes to the DDIM sampler or the probability flow ODE, the convergence analysis was originally provided in Chen et al., 2023c , with a set of subsequent work devoted to sharpening the iteration complexity (Li et al., 2024c, ; Huang et al., 2024a, ; Li et al., 2024d, ; Liang et al.,, 2025). In addition, a recent line of work Li and Yan, 2024a ; Huang et al., 2024b ; Azangulov et al., (2024); Potaptchik et al., (2024); Liang et al., (2025); Tang and Yan, (2025) explored how DDPM and DDIM adapt to unknown low-dimensional structures underlying the target data distribution, leading to substantially improved theoretical guarantees. Finally, rather than reducing the number of score evaluations, a strand of recent work (e.g., Shih et al., (2023)) accelerates diffusion sampling by running denoising steps in parallel. Interestingly, our proposed method is also amenable to parallel computing to some extent, although this is beyond the main goal of the current paper.
@@ 2 Preliminaries
where $X_{0}$ and $Z$ are independently generated. Clearly, this construction obeys
@@
In the sequel, we shall often refer to $\tau$ as the time variable too, as long as it is clear from the context. We will develop and describe our algorithm largely based on this continuous-time forward process (4).
@@
for any $x\in\mathbb{R}^{d}$, which readily implies that
@@
As has been shown by Song et al., (2021), there exists an ordinary differential equation (ODE), called the probability flow ODE, that is able to reverse the forward process (4) (in the sense of yielding matching marginal distributions). Note that the precise form of the probability flow ODE depends on the parameterization of the forward process (e.g., $(X_{t})$ and $(\overline{X}_{\tau})$ are associated with different, although closely related, ODEs). In particular, the probability flow ODE associated with process (4) takes the following form:
@@ 3 Algorithm
In this section, we present the main ideas and detailed procedure of the proposed algorithm. At a high level, our algorithm is an iterative procedure that attempts to approximately solve ODE (9) by approximating the true scores in the integrand with the aid of high-order polynomials.
@@
Recall the fact that $X_{t}=\overline{X}_{1-\overline{\alpha}_{t}}$ and $X_{t-1}=\overline{X}_{1-\overline{\alpha}_{t-1}},$ which makes apparent the importance of time points $1-\overline{\alpha}_{t}$ and $1-\overline{\alpha}_{t-1}$ when generating $Y_{t-1}$. Armed with the probability flow ODE (9), an ideal strategy for step t1t-1 would be to evaluate the following integral:
@@
However, computing (10) involves evaluation of an infinite number of score functions (i.e., those in the integral), making it computationally intensive. As a result, everything comes down to approximating the integral using a small number of score function evaluations.
1 input: $\tau_{t,0}=\tau_{t+1,K-1}=1-\overline{\alpha}_{t}$ for all $1\leq t\leq T-1$, and $\tau_{T,0}=1-\overline{\alpha}_{T}$. Choose $\tau_{t,i}$ as in (11).
2 initialization: $x_{\tau_{T,0}}\leftarrow Y_{T}\sim\mathcal{N}(0,I_{d})$.
3 for $t=T,\dots,2$ do
@@ 7 $Y_{t-1}\leftarrow x_{\tau_{t,K-1}}^{(N)}$
High-order polynomials.
In comparison to first-order solvers like DDIM that construct $Y_{t-1}$ solely based on the score estimate $s_{1-\overline{\alpha}_{t}}(\cdot)$, we propose to take advantage of the score functions at $K$ time points — denoted by $\tau_{t,0},\dots,\tau_{t,K-1}$ — within the interval $1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$. To be precise, set
@@
to be the two endpoints of $1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$, and create equi-spaced points within this interval as
@@
$\psi_{i}(\tau)\coloneqq\frac{\prod_{i^{\prime}:i^{\prime}\neq i}(\tau-\tau_{t,i^{\prime}})}{\prod_{i^{\prime}:i^{\prime}\neq i}(\tau_{t,i}-\tau_{t,i^{\prime}})},\qquad 0\leq i<K.$ (12)
Given a set of points $(\tau_{t,i},(1-\tau_{t,i})^{-3/2}s_{\tau_{t,i}}(Y^{\mathsf{ode}}_{\tau_{t,i}}))$ for $0\leq i<K$, we would like to approximate $\frac{1}{(1-\tau)^{3/2}}s^{\star}_{\tau}(x_{\tau})$ via the Lagrange interpolating polynomial passing through the above $K$ points, i.e.,
@@
Here, we have also replaced the true score $s_{t}^{\star}$ with the score estimate $s_{t}$ on the right-hand side of (13). Note that the Lagrange interpolating polynomial on right-hand side of (13) forms a unique degree-(K1)(K-1) polynomial passing through these $K$ points, and we take advantage of $K$ score function evaluations to approximate $\frac{1}{(1-\tau)^{3/2}}s_{\tau}^{\star}(Y^{\mathsf{ode}}_{\tau})$.
@@
Noteworthily, the polynomial appproximation (13) still cannot be readily used, given that the points $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$ are inaccessible in general. To remedy this issue, we propose a successive refinement scheme that alternates between estimating $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$ and computing the the associated scores. More precisely, let us perform an iterative estimation procedure containing $N$ rounds, where in each round $n$, we produce a sequence $\{x_{\tau_{t,i}}^{(n)}\}_{0\leq i<K}$ as an estimate of $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$. When $\{x_{\tau_{t,i}}^{(n)}\}_{0\leq i<K}$ is available, we compute, for each $\tau^{\prime}\in\{\tau_{t,0},\dots,\tau_{t,K-1}\}$,
@@
In a nutshell, we employ $\{x_{\tau_{t,i}}^{(n)}\}_{0\leq i\leq K-1}$ to update the scores used in the polynomial approximation (13), which in turn helps us generate $\{x_{\tau_{t,i}}^{(n+1)}\}_{0\leq i\leq K-1}$ as inspired by the ODE (9). This strategy is expected to yield improved estimates of $\{Y^{\mathsf{ode}}_{\tau_{t,i}}\}$ as the number of rounds increases.
@@
With the above ideas in mind, we are now ready to present our algorithm.
@@
  • Iterative update rule. Working backward from $t=T$ to $t=1$, we generate the point $Y_{t-1}$ as follows. Look at the time points $\{\tau_{t,i}\}_{0\leq i<K}$, and initialize both the sequence $\{x_{\tau_{t,i}}^{(0)}\}_{0\leq i\leq K-1}$ and $x_{\tau_{t,0}}$ using the value $x_{\tau_{t+1,K-1}}^{(N)}$ obtained in the previous iteration. For each round $n=0,\dots,N-1$, we generate a sequence $\{x_{\tau_{t,i}}^{(n+1)}\}_{0\leq i\leq K-1}$ according to Eq. (14). After $N$ rounds of successive refinement, we output $Y_{t-1}=x_{\tau_{t,K-1}}^{(N)}$.
@@
Several remarks are in order. To begin with, once the value of $Y_{T}$ is given, the sequence $\{x_{\tau_{t,i}}^{(n)}\}$ is purely deterministic. Thus, when accurate score functions are available at the requested time points, the discrepancy between the distributions of $Y_{t}$ and $X_{t}$ mainly stems from the approximation error when interpolating a complicated function using polynomials of degree K1K-1. In addition, our algorithm requires computing the score functions at the sequence $\{x_{\tau_{t,i}}^{(n)}\}_{0\leq i\leq K-1,0\leq n\leq N,1\leq t\leq T}$, and hence the total number of score function evaluations is $TKN$. As we shall see momentarily in our theoretical development, $K$ is taken to be a constant, whereas $N$ is chosen to be logarithmically large. Consequently, the total number of score function evaluations is on the order of $\widetilde{O}(T)$.
Remark 1 (The time indices $t$ vs. $\tau$).
In Algorithm 1, we employ both the discrete-time index $t$ and the continuous variable $\tau$, both of which are instrumental in describing our algorithm. Recall that $t$ is a discrete step-index for both the forward and backward processes in discrete time, while $\tau$ is the continuous time variable (cf. (9)) taking value in $[0,1]$. At Step $t$, we iteratively refine our estimate of $Y_{t-1}$, or equivalently, $Y_{1-\overline{\alpha}_{t-1}}^{\sf ode}$, at $K$ time points within the interval $1-\overline{\alpha}_{t-1},1-\overline{\alpha}_{t}]$, where these intermediate points do not need to coincide with the time indices of the actual discrete-time process.